OpenStudy (anonymous):

8.1 is what % of 162

OpenStudy (anonymous):

\(x \%\) means the fraction \(\frac{x}{100}\) of something. Try solving this then:\[8.1=\left(\frac{x}{100}\right)\times 162\]

OpenStudy (anonymous):

 Now what?